# RollingCache: Using Runtime Behavior to Defend Against Cache Side Channel Attacks

### v1

@@ 1 Introduction
We present RollingCache, a cache design that utilizes the dynamic nature of the runtime state of the system to create the necessary nondeterminism in the addresses that contend in cache. RollingCache uses indirections for mapping a set of addresses to cache sets. These indirections are not static for a predetermined period, and get updated to a new location after a certain number of cache line fills. Rather than a replacement in the same cache set when the fill limit is reached, a conflict miss updates the indirection, without requiring data relocation. The rate at which the mapping changes is a nondeterministic function of the number of replacements in the cache.
@@ 5 Security Evaluation
  • Unknown number of cache fills available
@@
We describe how these features render an attack non-conclusive in the following sections.
5.1 Prime Sequence: Toy Example Attack
@@
The probability that an individual access remains in the allocated CacheSet after priming can be expressed using the following equations, where the probability of scenario ‘A”, i.e., prime begins at a pointer rollover is expressed as P(A), P(B) is the probability of scenario ‘B’ (rollover during priming), and P($C_{n}$) is the probability of no random eviction of the $n$th access in the prime sequence by any of the following accesses in the sequence, and ‘W’ is the number of ways in a cacheSet.
@@
$P(1\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ a\ ptr\_rollover)and\\ P(no\ random\_eviction)\\ =P(A)*P(C)\\ =(1/W)(\frac{W-1}{W})^{W-1}$ (4)
@@
$P(2\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ ptr\_rollover)or\\ P(primedoesn^{\prime}t\ start\ at\ ptr\_rollover)*P(no\ random\_eviction)\\ =P(A)+P(B)*P(C)\\ =(1/W)+((W-1)/W)(\frac{W-1}{W})^{W-1}\\ =(1/W)+(\frac{W-1}{W})^{W}$ (5)
$P(3\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ ptr\_rollover)*P(no\ random\_eviction)\\ or\ P(prime\ doesn^{\prime}t\ start\ at\ ptr\_rollover)(1)\\ =P(A)P(C)+P(B)\\ =(1/W)(\frac{W-1}{W})^{W-1}+((W-1)/W)$ (6)
@@
The data left in the CacheSets after the prime depend on the original state of the cache. Incurring a miss at probe itself can trigger another pointer update, and can undo the effect of priming, which implies that the attacker may want to infer a victim’s access behavior based on the very first miss at probe. However, as illustrated in Figure 5, there is no good candidate for probe, as all of 1, 2, 3 can be missing even before the victim’s access due to random eviction or due to pointer rollover. 2, 3, 4 are also not good candidates for probe since a subsequent hit on these does not imply that the victim did not access the same AddrSet. They leave the possibility of the victim going undetected as shown above in Figure 5,‘B’.
Figure 5: A toy example attack sequence under RollingCache
@@ 5.2 Probe Inference
We determine the probability of a victim causing a miss to one of the accesses that were primed, and then find the likelihood of those misses being due to the victim accessing the addrSet of interest. Let us consider a RollingCache with ‘S’ CacheSets. The number of distinct AddrSets is also ‘S’. Let us say that the attacker tries to determine if the victim accesses a specific AddrSet $AddrS_{i}$. The first phase of the attack constitutes making 2W accesses to addresses from AddrSet $AddrS_{i}$. The attacker then allows the victim to make its access, and probes one of the addresses from the prime sequence.
To determine the right candidate for probe, we evaluate individual probabilities of accesses resulting in a miss if the victim did indeed access the AddrSet of interest $AddrS_{i}$.
$P(Miss1\ due\ to\ AddrSi)=P(A)P(C)(1)\\ =(1/W)((W-1)/W)^{W-1}$ (8)
Since 2 can be evicted by the victim accessing AddrSi only if the prime had started at pointer rollover
$P(Miss2\ due\ to\ AddrSi)=(1/W)$ (9)
@@
$P(Miss3\ due\ to\ AddrSi)=0$ (10)
$P(Miss4\ due\ to\ AddrSi)\\ =P(not\ starting\ prime\ after\ a\ ptr\ rollover)\\ and\ P(random\_eviction\ due\ to\ victim)\\ =((W-1)/W)(1/W)$ (11)
From equations 8, 9, 10, 11, the probability of miss due to victim accessing the same AddrSet $AddrS_{i}$ is the highest for probing access ‘2’. This might be the access of interest to the attacker for probing victim’s access. We now calculate the probability of the victim’s access actually causing a miss on the access, which can be obtained using eq. 9 with individual probabilities of the victim’s access being from different addrSets.
$P(Miss2\ due\ to\ victim)\\ =P(accessing\ AddrSi)*P(M2\ due\ to\ AddrSi)+\\ P(accessing\ any\ other\ AddrSj)*P(M2\ due\ to\ AddrSj)$ (12)
Since all the S sets are equally likely, the probability of the access ebing from $AddrS_{i}$ or $AddrS_{j}$ is 1/S and there are (S-1) other addrSets besides $AddrS_{i}$
$P(Miss2\ due\ to\ AddrSj)=(P(A)+P(B)*P(C))\frac{1}{SW}\\ =[\frac{1}{W}+\frac{W-1}{W}*(\frac{W-1}{W})^{W-1}](\frac{1}{S})(\frac{1}{W})$ (13)
Using 9, 13 in 12
$P(Miss2\ due\ to\ victim)\\ =(\frac{1}{SW})+[\frac{1}{W}+\frac{W-1}{W}*(\frac{W-1}{W})^{W-1}](\frac{1}{SW})(\frac{S-1}{S})\\ =\frac{1}{SW}[1+(\frac{S-1}{S})(\frac{1}{W}+(\frac{W-1}{W})^{W})]$ (14)
For S=64 and W=2, the above probability of getting a miss due to the victim’s access is 0.0135, i.e the probability of getting a miss on 2 due to a victim access is 1.35%, which could still be an access from the same AddrSet $AddrS_{i}$ or another random AddrSet.
[…]
@@
$P(A_{i}|B)=\frac{P(B|A_{i})P(A_{i})}{\sum\limits_{j}P(B|A_{j})P(A_{j})}$ (15)
[…]
$P(AddrS_{i}|M)=\frac{P(M|AddrS_{i})P(AddrS_{i})}{\sum\limits_{AddrS_{j}=1}^{S}P(M|AddrS_{j})P(AddrS_{j})}$ (16)
@@
As seen above in the previous subsection, the probability of miss due to victim’s access to the same AddrSet is maximum in eq.9
$P(M|AddrS_{i})=(1/W).$ (19)
The probability of incurring a miss due to an access from some other AddrSet, $AddrS_{j}$, can be given as the probability of CacheSet PsPtr[$AddrS_{i}$] being occupied by $AddrS_{j}$. Since there are $S$ CacheSets to choose from, each of these probabilities is $1/S$.
The attacker’s probe access resulting in a miss is likely to coincide with the location of a victim’s access 1/W times. Assuming the attacker has the probe resolution of timing a single access by the victim, the
$P(M|AddrS_{j})=\\ (1/W+((W-1)/W)*((W-1)/W)^{(W-1)})(1/S)(1/W)$ (20)
Using 17181920 in 16 we get,
$P(AddrS_{i}|M)=\\ \frac{(\frac{1}{W})(\frac{1}{S})}{(\frac{1}{W})(\frac{1}{S})+(S-1)(\frac{1}{S})(\frac{1}{W}+(\frac{W-1}{W})(\frac{W-1}{W})^{(W-1)})(\frac{1}{SW})}$ (21)
We can calculate this probability for different number of CacheSets. Using ‘S’ = 64 and ‘W’ = 2:
$P(AddrS_{i}|M)=0.5752$ (22)
The analysis in essence estimates the chances that accesses used to prime the cache actually reside in the cache; and the victim’s access causes a miss for probe on that access; and the miss is because of the access to the same address set as the one being primed. The probability of getting a miss due to victim accessing $AddrS_{i}$ is 57.52% of 1.35%, i.e 0.77%. The attack requires successful priming of the cacheSet and eviction due to victim accessing the same addrSet, this analysis shows that the likelihood of all these events happening under the design of RollingCache is 0.77%.
If we consider no attacker control over the victim’s accesses, the introduced error rates make iterations of attack to decipher the victim’s access ineffective. Further, repeating the attack will not see the same behaviour as the contention for a second run of the attack might be in other cacheSets with different states from the first run.
@@ 5.4 AddrSet Cache Footprint
5.5 Cache Flush
In the case of a cache flush instruction, as the address being flushed may share a CacheSet with different AddrSets and might not reside in a specific CacheSet, the attacker cannot get control of complete CacheSets using cache flush instructions. Further, RollingCache retains the state of the indirection pointers (PrPtr, PsPtr, and all associated metadata) across flushes. When the victim resumes execution, its accesses would continue to be directed to the CacheSet pointed to by the PrPtr of the accessed AddrSet but the number of relocations in the CacheSet is not impacted by the flushes. The attacker would not be able to decipher the AddrSet that was accessed by the victim if it incurs a miss in some CacheSet.
5.6 Robustness against replay attacks
It has been shown that finding an eviction set for Ceaser-S within the remap interval is possible, because there is a static hashmap within the remap window. Scattercache similarly has a constant set of partially congruent addresses, which once discovered can be reused for launching attacks [26]. Unlike both Scattercache and Ceaser-S, RollingCache does not have specific address sets contending with each other for a specified duration, rather the mapping updates every W replacements, and we support a maximum of 2W different address lookups. This absence of an epoch is one crucial difference from prior works of defense, blocking discovery or reuse of set of addresses contending in RollingCache. The paper attacking Randomized caches uses clflush [26] to reuse the discovered partially congruent eviction sets for priming. Clflush cannot be used to prime a CacheSet in Rolling Cache, since we do not write to the same freed lines in a CacheSet. The addresses being flushed could be mapped to the CacheSet pointed to by the past or the present pointer. The writes after flush continue to contend with yet another address set, since the mapping updates every W replacements. The congruent or partially congruent addresses cannot be reused, as the mapping does not last beyond W accesses.
Further, the attacks on random replacement without filling the cache rely on shared data reuse and the clflush instruction, which is a different threat model compared to contention-based attacks. This work is a defense against contention based attacks, not against reuse based attacks, and is complementary to defense systems built to handle reuse based attacks, such as TimeCache [16].
RollingCache uses random replacement, and the attack RELOAD+REFRESH [4] discovers and exploits LRU and other non-random replacement policies. The paper demonstrates attacks that can discover and exploit LRU, Tree PLRU, CLOCK, NRU, Static and Bimodal RRIP, and states that the absence of randomness makes the attacks possible. Given that RollingCache uses random replacement, the attack will not be effective. Casa is a framework for analyzing random caches and their epochs for their vulnerability and discovery of eviction sets using repeated accesses [3]. Since RollingCache has no notion of epoch, does not allow repeated contention on the same address, or in other words does not have a static mapping for a short duration or epoch, the technique in CaSa does not apply to the eviction set inference in RollingCache.

### v2

@@ Abstract
§§footnotetext: Work done while at University of Rochester.
@@ 1 Introduction
We present RollingCache [14], a cache design that utilizes the dynamic nature of the runtime state of the system to create the necessary nondeterminism in the addresses that contend in cache. RollingCache uses indirections for mapping a set of addresses to cache sets. These indirections are not static for a predetermined period, and get updated to a new location after a certain number of cache line fills. Rather than a replacement in the same cache set when the fill limit is reached, a conflict miss updates the indirection, without requiring data relocation. The rate at which the mapping changes is a nondeterministic function of the number of replacements in the cache.
@@ 5 Security Evaluation
  • Unknown number of cache line fills available prior to a rollover
@@
We describe how these features render an attack inconclusive in the following sections.
5.1 Priming Sequence
@@
The probability that an individual access remains in the allocated CacheSet after priming can be expressed using the following equations, where the probability of scenario ‘A’, i.e., prime begins at a pointer rollover is expressed as P(A), P(B) is the probability of scenario ‘B’ (rollover during priming), and P($C_{n}$) is the probability of no random eviction of the $n$th access to a set, by any of the following accesses in the sequence, and ‘W’ is the number of ways in a CacheSet.
@@
$P(1\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ a\ ptr\_rollover)and\\ P(no\ random\_eviction)\\ =P(A)*P(C_{1})\\ =(1/W)(\frac{W-1}{W})^{W-1}$ (4)
@@
$P(2\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ ptr\_rollover)or\\ P(primedoesn^{\prime}t\ start\ at\ ptr\_rollover)*P(no\ random\_eviction)\\ =P(A)+P(B)*P(C_{1})\\ =(1/W)+((W-1)/W)(\frac{W-1}{W})^{W-1}\\ =(1/W)+(\frac{W-1}{W})^{W}$ (5)
$P(3\ in\ cache\ after\ Prime)\\ =P(prime\ starts\ at\ ptr\_rollover)*P(no\ random\_eviction)\\ or\ P(prime\ doesn^{\prime}t\ start\ at\ ptr\_rollover)(1)\\ =P(A)P(C_{1})+P(B)\\ =(1/W)(\frac{W-1}{W})^{W-1}+((W-1)/W)$ (6)
@@
The data left in the CacheSets after the prime depend on the original state of the cache. The probability of prime access 1, 2, 3, and 4 remaining in the cache at the end of the prime sequence is 0.25, 0.75, 0.75, and 1 respectively. The corresponding probability of eviction during priming and hence a miss for the probe of access $x$ due to prime ($MPrime_{x}$) is 0.75, 0.25, 0.25, and 0.
Incurring a miss at probe itself can trigger another pointer update, and can undo the effect of priming, which implies that the attacker may want to infer a victim’s access behavior based on the very first miss at probe. However, as illustrated in Figure 5, there is no good candidate for probe, as all of 1, 2, 3 can be missing even before the victim’s access due to random eviction or due to pointer rollover. 2, 3, 4 are also not good candidates for probe since a subsequent hit on these does not imply that the victim did not access the same AddrSet. They leave the possibility of the victim going undetected as shown above in Figure 5,‘B’.
Figure 5: Example attack sequence for a 2-way RollingCache
@@ 5.2 Probe Inference
Let us consider a RollingCache with ‘S’ CacheSets. The number of distinct AddrSets is also ‘S’. Let us say that the attacker tries to determine if the victim accesses a specific AddrSet $AddrS_{i}$. The first phase of the attack constitutes making 2W accesses to addresses from AddrSet $AddrS_{i}$. The attacker then allows the victim to make its access, and probes one of the addresses from the prime sequence.
We evaluate the individual probabilities of probe accesses to each possible primed address resulting in a miss if the victim did indeed access the AddrSet of interest $AddrS_{i}$.
$P(Miss1\ due\ to\ AddrS_{i})=P(A)P(C_{1})(1)\\ =(1/W)((W-1)/W)^{W-1}$ (8)
Since 2 can be evicted by the victim accessing $AddrS_{i}$ only if the prime had started at pointer rollover
$P(Miss2\ due\ to\ AddrS_{i})=P(A)=(1/W)$ (9)
@@
$P(Miss3\ due\ to\ AddrS_{i})=0$ (10)
$P(Miss4\ due\ to\ AddrS_{i})\\ =P(B)*P(random\_eviction\ due\ to\ victim)\\ =((W-1)/W)(1/W)$ (11)
For W=2, the above equation 8, 9, 10 and 11 i.e the probability of incurring a miss on a probe of prime access 1, 2, 3, and 4, given a victim access to $AddrS_{i}$, is 0.25, 0.5, 0, and 0.25 respectively.
The probability of incurring a miss on a probe of access $x$ due to a victim’s access of some other AddrSet, $AddrS_{j}$, depends on the probability of $x$ remaining in the cache after prime, and the probability of the CacheSet being written by $AddrS_{j}$. Since there are $S$ CacheSets, the probability is $1/S$ assuming a uniform and random mapping distribution. The attacker’s probe access resulting in a miss is likely to coincide with the location of a victim’s access with a probability 1/W.
$P(Miss1\ due\ to\ AddrS_{j})=(P(1incacheafterPrime)\frac{1}{SW}\\ =[\frac{1}{W}*(\frac{W-1}{W})^{W-1}]\frac{1}{SW}$ (12)
$P(Miss2\ due\ to\ AddrS_{j})=(P(2incacheafterPrime))\frac{1}{SW}\\ =[\frac{1}{W}+\frac{W-1}{W}*(\frac{W-1}{W})^{W-1}]\frac{1}{SW}$ (13)
$P(Miss3\ due\ to\ AddrS_{j})=(P(3incacheafterPrime))\frac{1}{SW}\\ =[\frac{1}{W}*(\frac{W-1}{W})^{W-1}+\frac{W-1}{W}]\frac{1}{SW}$ (14)
$P(Miss4\ due\ to\ AddrS_{j})=(P(4incacheafterPrime))\frac{1}{SW}\\ =\frac{1}{SW}$ (15)
We can calculate the probability of a probe incurring a miss on the $x$th access in the prime sequence due to the victim’s access to any AddrSet in the cache as follows:
$P(M_{x}\ due\ to\ victim)\\ =P(accessing\ AddrS_{i})*P(M_{x}\ due\ to\ AddrS_{i})+\\ P(accessing\ any\ other\ AddrS_{j})*P(M_{x}\ due\ to\ AddrS_{j})$ (16)
Since all the S sets are equally likely, the probability of the access being from $AddrS_{i}$ or $AddrS_{j}$ is 1/S and there are (S-1) other AddrSets besides $AddrS_{i}$.
Using equations 12, 13, 14, 15 and for S=64 (and W=2), the probability of getting a miss due to the victim’s access to any AddrSet in equation16 is 0.0058, 0.0136, 0.0058, and 0.0116 respectively on a probe access to the first, second, third, and fourth prime accesses respectively.
To find:  The probability of the victim’s access being from AddrSet $AddrS_{i}$, given a miss is incurred by the attacker during probe, i.e., P($AddrS_{i}|M_{x}$).
@@
$P(A_{i}|B)=\frac{P(B|A_{i})P(A_{i})}{P(B)}$ (17)
@@
Using 17, we can express the P($AddrS_{i}|M_{x}$) in terms of the probability of a miss occurring due to the victim accessing different AddrSets or due to an eviction during priming (miss due to prime ($MPrime_{x}$)) as follows:
$P(\frac{AddrS_{i}}{M_{x}})=\frac{P(\frac{M_{x}}{AddrS_{i}})P(AddrS_{i})}{P(MPrime_{x})+P(M_{x}\ due\ to\ victim)}$ (20)
  • For S=64 (and W=2), the probability of a miss due to the victim having accessed $AddrS_{i}$ given a miss on a probe of the first, second, third, and fourth prime accesses is 0.0052, 0.0299, 0, and 0.3368 respectively. This probability is expected to be 1 in conventional cache.
  • The corresponding probabilities of miss due to victim having accessed a different AddrSet given a miss at probe are similar, i.e 0.0025, 0.0217, 0.0227, 0.6632 respectively. This probability for a conventional cache is expected to be 0. The probability is low for the first three prime accesses because of the high probability of self-eviction at prime ($P(MPrime_{x})$ is high). Hence, this makes it hard for a probe of the third access to infer that the miss was due to the victim.
  • Additionally, based on P($Miss_{x}\ due\ to\ AddrS_{i}$), 75%, 50%, 100%, and 75% of the time, the access to $AddrS_{i}$ by the victim will go undetected due to a lack of a miss on the corresponding probe access, further reducing the probability of information leak for all probe accesses, including that for a probe of the fourth prime access.
In summary, RollingCache’s design introduces nondeterministic behavior at the prime, victim, and probe accesses, resulting in high error rates in victim access prediction and making iterations of attack to decipher the victim’s access ineffective. Unlike in a conventional cache, inferring the AddrSet accessed from the misses is not deterministic due to contention spread across different AddrSets. Further, repeating the attack will not see the same behavior, as the contention for a second run of the attack might be in other CacheSets with different states from the first run.